### v2

Abstract
In this document, we discuss a multi-step approach to automated construction of a knowledge graph, for structuring and representing domain-specific knowledge from large document corpora. We apply our method to build the first knowledge graph of nuclear fusion energy, a highly specialized field characterized by vast scope and heterogeneity. This is an ideal benchmark to test the key features of our pipeline, including automatic named entity recognition and entity resolution. We show how pre-trained large language models can be used to address these challenges and we evaluate their performance against Zipf’s law, which characterizes human-generated natural language. Additionally, we develop a knowledge-graph retrieval-augmented generation system that combines large language models with a multi-prompt approach. This system provides contextually relevant answers to natural-language queries, including complex multi-hop questions that require reasoning across interconnected entities.
Index Terms:
Retrieval Augmented Generation, Knowledge Graph, Nuclear Fusion, Information Extraction, Large Language Models.
I Introduction
Nuclear fusion has the potential to transform the global energy landscape by providing a sustainable source of carbon-free energy to our society. However, the achievement of nuclear fusion power presents numerous challenges that extend beyond the inherent complexities of nuclear physics and require multidisciplinary approaches, e.g., [1, 2].
As the area of research continues to expand, the need for an efficient system of data management and knowledge retrieval has become increasingly critical [3, 4, 5]. Addressing this challenge requires not only improving data accessibility but also enabling seamless interaction between human reasoning and machine processing. An interconnected system for online data, that can bridge this gap, holds significant appeal to many disciplines beyond nuclear fusion energy, as it has been clearly illustrated in [6, 7]. To address this need, new data models have been developed based on graph structures [8], and ontology frameworks [9].
In a graph, knowledge is stored in nodes (entities) linked together by edges. Nodes encapsulate concepts, objects, or nouns relevant to the domain, while edges define relationships between them. These relationships can carry semantic meaning like in subject-relationship-object triplets, or represent simpler connections such as the co-occurrence of two entities within the same text.
Ontology models, on the other hand, define a vocabulary of terms, their meanings, and organize them in hierarchical classes connected by well defined relationships. These two approaches represent complementary data models that, when used together, can enhance the accuracy and richness of the final Knowledge Base (KB), e.g., [10].
The development of a KB of fusion energy, which organizes information into structured categories and representations, will accelerate the development of more ”FAIR” data sharing, disseminating and referencing [3, 4]. Over time, such a system will foster collaborations between all partners in a fusion supply chain, helping to de-risk the design, development and construction of fusion power plants, while also reducing the barriers to understanding between field experts, funding agencies and policy makers.
In this work, we present an automated approach to generate a graph-based KB or Knowledge Graph (KG) of nuclear fusion energy starting from a large corpora of scientific documents and by leveraging the inference power of pre-trained Large Language Models (LLMs). Additionally, we developed a KG Retrieval-Augmented Generation (KG-RAG) machine. This combines the advanced linguistic capabilities of modern generative AI with domain-specific knowledge derived from previously unseen sources. The KG-RAG enables the effective retrieval of domain-specific information and the generation of accurate answers to user queries helping minimizing model hallucinations, i.e., the production of plausible but fictitious responses.
II Related works
The creation of a KB begins with the process of Information Extraction (IE) from a text corpus. To automate the learning of linguistic features, different trainable algorithms have been developed over the years. These include approaches based on statistical analysis [11, 12], Bayesian inference [13, 14], and machine learning algorithms, e.g., [15, 16]. Early methods often relied on precompiled heuristic rules and human supervision to extract linguistic patterns e.g., [11] as well as on hand-annotated training sentences [17]. Other models employed ontologies to define classes and relationships of interest along with hand-labelled training texts, e.g., [15, 18]. For these kinds of Natural Language Processing (NLP) tasks, manual efforts remain a significant bottleneck, especially when dealing with large text corpora.
An essential task in IE is Named Entity Recognition (NER), which identifies entities specific to a particular domain within a given text. This process is often followed by Relation Extraction (RE), which establishes relationships between entities. Recurrent neural networks have been extensively used to accomplish these tasks, e.g., [19] and [20]. However, recent advancements in AI-powered NLP, particularly with LLMs, have enabled machines to assimilate vast amounts of text and process it with capabilities once believed to be uniquely human. Pre-trained LLMs have demonstrated good IE performance across disciplines, in both zero- and few-shot scenarios [21, 22, 23]. While their performance tends to decline when handling complex tasks [24] it can be recovered by breaking down high-level objectives into simpler, more manageable sub-tasks, each addressed through targeted prompts[25, 26].
The significant advancements, made over the years in the field of IE, have allowed the creation of several cross-disciplinary online KBs. Some of which were manually curated by their respective communities, as in the case of Freebase KG. Others, like DBpedia, were generated automatically from web sources such as Wikipedia [27]. To the best of our knowledge, a KB dedicated to nuclear fusion energy is not included in any of these databases, despite its societal relevance and the increasing interest shown by policymakers and the public. To address this gap, we embarked on the creation of the first KG for nuclear fusion energy.
Figure 1: Zipf’s law applied to the top-ranked 500 single word entities in a case study of 349 abstracts: (a) before entity resolution, (b) after entity resolution process. The legend shows the Zipf parameter $(C)$ extracted from the fit and the normalized $\chi^{2}$ value.
III Methods
Our work can be divided in two main parts: the automated construction of a KG and the development of a KG-RAG. These are discussed in the next two sections.
III-A Automated construction of a KG
TABLE I: Data scoping
Key Value
Summary:
Abstracts 8358
Total words 7799769
Time range 1958-2024
Search patterns:
Tokamak 4506
Stellarator 1632
Inertial confinement 1202
Nuclear fusion 328
Fusion energy 355
Deuterium-Tritium 335
A schematic of the pipeline for the automated construction of a KG is shown in figure 2. The workflow consists in the following steps: Data Acquisition (DAQ), NER, entity resolution, KG construction and RE.
Figure 2: Workflow for the automated creation of a KG. The first layer consists in the data acquisition. This involves scoping, to identify relevant data sources, formatting, to structure the data in JSON format. The second layer involves NER. This step is achieved through prompt-driven processing with a LLM. Following NER, a second LLM-assisted step performs entity resolution. This generates a set of rules for standardizing entities. The following step integrates the NER output and resolution rules to construct the KG. Entities that frequently appear together in the same context are paired. For each pair, the LLM extracts a relationship linking the two entities.
The DAQ layer involves two steps:
  • Scoping. This step identifies relevant data sources, which is accomplished by using an API to access the online archive of scholarly publications at www.lens.org.
  • Formatting. In this step, the retrieved data is structured in JSON format.
We downloaded 8358 scientific abstracts from journal articles related to nuclear fusion energy and associated metadata such as: authors, keywords, publication year and scholarly citation count. The search process involved looking up specific patterns in article titles. A summary of the dataset is provided in Table I, along with the list of search patterns used. This approach may include a small number of articles that are not strictly related to nuclear fusion energy but match the search criteria1118 out of 335 in the Deuterium-Tritium sample. To filter out irrelevant abstracts, we implemented a data validation step using an LLM and a keyword-based filter. In this study, we have used Llama3.1 or LLama 3.3 by Meta, deployed on Amazon Web Services (AWS) equipped with 405 and 70 billion parameters, respectively 222meta.llama3-1-405b-instruct-v1:0 and meta.llama3-3-70b-instruct-v1:0.
Following the DAQ stage, the LLM was tasked with performing NER on the selected abstract text. To improve extraction accuracy, the text was split into individual sentences to reduce input length.
Previous works provided a definition of entity in their prompts effectively capturing the scope of their studies. Nevertheless, prompting for NER can be accomplished in few- or zero-shot mode depending on whether examples of entities are passed to the prompt or not, [21, 23, 25, 26]. In the present work, we do not provide explicit examples of entities but we do guide the extraction process by suggesting a list of category types to which entities belong. Although this method requires minimal user input, it allows to categorize entities in standard, cross-model communities (or groups) that may be used during the retrieval stage. To achieve optimal categorization, we have chosen category types that are as exclusive as possible, while also being broad enough to encompass a significant portion of relevant entities. Further refinement of these category types may be accomplished by running another NER pass with an updated LLM prompt. A list of the category types used in this work is given in the appendix. The accuracy of the classification process, that assigns each entity to a category type in the list, was estimated by comparing randomly selected entities and their LLM-assigned categories against the corresponding manually assigned ones. We acknowledge the limitations of this evaluation, as it depends on the specific choice of category types. Additionally, this estimate likely represents an upper bound on classification accuracy, given that some of the categories are not entirely exclusive and an entity may be correctly assigned to multiple categories. We found that the LLM correctly classified $\simeq$94of the entities  [^3] The accuracy was seen to decrease for LLMs with a smaller amount of parameters such as LLama3.1 8B.
Following the NER stage, the entity resolution step is performed which allows to resolve duplicates, acronyms and other inconsistencies among entities. During this process, entities are automatically standardized by using singular forms and lowercase spellings (when appropriate), while entities consisting solely of numbers are excluded. Additionally, this step removes entities inferred by the LLM that are not explicitly mentioned in the text. Part of entity resolution involves an additional LLM prompting pass. To this aim, a dictionary of top-ranked entities is created from the NER data and embedded into a prompt along with the simple instruction to identify acronyms with their expanded forms. Similar instructions are used to standardize chemical elements. The LLM then produces a list of substitutions required to transform entities into their standardized forms. This list is parsed during the graph creation stage ensuring consistency and reducing noise in the final knowledge graph.
The entity resolution effectiveness can be appreciated by looking at the plots in figure 1. These show the frequency of the 500 top-ranked single word entities fitted by the Zipf law, before and after de-noising stage.
The Zipf law is an empirical principle often observed in NLP and other areas of linguistics. It states that in a given corpus, the frequency $(f)$ of any word is inversely proportional to its rank $(r)$ in the frequency table: $f=C/r$. After entity resolution, the frequencies of some entities change resulting in a better fit to Zipf’s law.
For the KG construction we have used the Neo4j framework. Each resolved entity is represented by an unique node in the graph. The graph structure is summarized in table II and includes nodes for abstracts, first authors, publication years, and key words, as well as individual sentences and entities. These nodes are interconnected through a set of predefined relationships, also summarized in table II and illustrated in figure 3:
TABLE II: Nodes and relationships properties
Label Key Value
Abstract name Abstract title
text Abstract text
url URL
citationCount Number of citations
Sentence name Sentence text
embeddings Vector embedding
Entity/Person/
TimeReference/KeyWord name Entity name
types Entity category type
edges Number of edges
Label Start/end Value
HAS_FIRST_AUTHOR Abstract/Person
WAS_PUBLISHED_IN Abstract/TimeReference
HAS_KEYWORD Abstract/KeyWord
HAS_SENTENCE Abstract/Sentence
CONTAINS Sentence/Entity
CC Entity/Entity weight
text (optional)
Figure 3: Example of KG accordingly to the graph architecture used in this work. In this diagram, (Entity 1, RELATIONSHIP, Entity 2) form a triplet. Triplets have both [CC] and semantic relationships, see text for more details.
[…]
III-B KG-RAG
Figure 4: Workflow of the KG-RAG. The user input is a question processed by a LLM that generates a Cypher query. A keyword search through the KG retrieves relevant documents (sentences) and triplets. These go through a sorter-filter module that yields a list of top-ranked sentences and triplets. Finally, the LLM generates an accurate answer to the user’s query.
The automated search of the KG for retrieval and generation purposes is accomplished by using the pipeline illustrated in figure 4. In the first step, the user submits a question and specifies the number of documents to retrieve. A first pass of LLM extracts relevant entities from the user’s question and embeds them into a Cypher query. Cypher is a declarative query language specifically designed for working with graph databases. Figure 5 shows two examples of LLM-generated Cypher queries generated from user inputs expressed in natural language. We considered both single-hop and multi-hop queries that can be answered by using information from a single node or multiple interconnected nodes, respectively.
We perform full-text search over the entire corpus to identify sentences that match entities extracted from the user’s query by the LLM. The retrieved sentences are then ranked in descending order of semantic similarity to the original question by a dedicated sorter module. Similarity scores are computed using the scalar product between the embedded vector representations of each sentence and the query. Sentence embeddings were generated using the ll-mpnet-base-v2 model. This maps a sentence into a 768 dimensional dense vector space and can be used for tasks like clustering or semantic search  [^5] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. Finally, the top-k abstracts containing the most relevant sentences are selected and incorporated into the prompt provided to the LLM for answer generation.
The retriever also searches the graph for triplets, i.e., subject-relationship-object. These are then filtered based on two criteria: they must contain entities recognized by the LLM in the user’s query or they must include verbs with matching lemma to that in the query. If no match is found, the system sorts all the triplets by the total occurrence of the subject-object entities. Finally, triplets are passed to the LLM for answer generation.
IV Results
In this study, we accomplished the automated construction of a KG of nuclear fusion energy that comprehensively represents the domain knowledge contained in 8358 scientific abstracts. To the best of our knowledge, this is the first KG in the domain of nuclear fusion energy. The graph structure is summarized in tables II and contains 108811 nodes and 718335 links.  [^6] The KG is available upon request: andrea.loreti@stfc.ac.uk
Assessing the performance of a RAG (or KG-RAG) system is complex, as it requires evaluating both retrieval accuracy and the faithfulness and correctness of generated responses, e.g., [28] and [29]. The interplay between RAG components, and the potential influence of the KG, make an exhaustive study of RAG performances beyond the scope of this work. However, we conducted two evaluation experiments as follows. First, we randomly selected $10$ abstracts and used ChatGPT to generate a specific question for each, simulating queries from field experts. We then examined the KG-RAG output to determine whether the corresponding abstracts were among the retrieved sources. In this experiment, the correct abstract was consistently ranked as the top result.
Figure 5: Two examples of LLM-generated Cypher queries, single-hop (left) and multi-hop (right), derived from user queries expressed in natural language. For generating Cypher queries, LLM was provided via prompt with the graph structure in table II.
Next, we generated a set of $50$ questions using a system prompt designed to simulate queries from different kinds of personas  [^7] We used Ragas: https://docs.ragas.io/en/v0.1.21/index.html. These questions were less specific, typically shorter, and contained typos or colloquial phrases. The retrieval accuracy for this broader set was $\simeq 50\%$. However, in all but one case, the retrieved documents remained highly relevant to the corresponding queries.
It must be noted that for all generated answers, a list of correct references is provided which links to the original text sources via URL. This result addresses a well-known problem associated with LLMs, specifically the issue of referencing the text source used in answer generation. In fact, even top-performing models currently available have been found to provide unsupported statements or incorrect, conflicting citations [30].
For what is concerned with the generative process, this primarily consists in summarizing the retrieved top-ranked documents and does not add extra information from knowledge acquired during the pre-training, unless explicitly prompted to do so. If, on the one hand, these guardrails may limit the creative scope of AI tools, they also help to prevent LLMs from hallucinating and producing fictitious contextual answers, thus addressing a well-known issue affecting LLMs, e.g., [31]. While this answer generation can be very accurate, it lacks abstraction capabilities. For instance, the tasks of summarizing the content of the whole KG or extracting information within a specific time range require mutli-hop capabilities that retrieve information beyond the content of the documents themselves. KG-RAG offers a powerful way to achieve this level of abstraction by leveraging the entities and relationships stored in the graph. For example, by utilizing the predefined relationships outlined in table II and the LLM-generated semantic relationships, it has been possible to provide KG-RAG with an initial layer of abstraction, as it can be seen in examples 2 and 3 below.
  • Example 1
    User: ”How does the presence of high-density carbon impurities affect the ignition and burn of deuterium-tritium (DT) plasma, and what are the implications for inertial confinement fusion (ICF)?”
    System: ”The presence of high-density carbon impurities has a negative effect on the ignition and burn of deuterium-tritium (DT) plasma in inertial confinement fusion (ICF) targets…”
    Sources:
    Abstract 1 [32]. One of the main concerns about the ignition and burn of deuterium-tritium DT plasma is the negative effect of impurities that can get into the thermonuclear fuel during target compression on the ignition of an inertial confinement fusion ICF target…
    Abstract 2 [33]. Accurate knowledge of the equation of state EOS of deuterium-tritium DT mixtures is critically important for inertial confinement fusion ICF. Although the study of EOS is an old topic, there is a longstanding lack of global accurate EOS data for DT within a unified theoretical framework… In the present work, we perform extensive state-of-the-art ab initio quantum Langevin molecular dynamics simulations to obtain EOS data for DT mixtures at densities from 0.1 gcm3 to 2000 gcm3 and temperatures from 500 K to 2000 eV, which are relevant to ICF processes…
    Abstract 3 [34]: Knowledge of the dynamics and concentration of impurities in the plasma core is crucial to the development of a deuterium-tritium DT tokamak fusion reactor…
  • Example 2
    User: ”Provide a list of the 3 most cited papers, in the database, published between 1958 and 2024; add the citation counts.”
    System: Based on the text, here is a list of the top 3 most cited papers along with their citation counts:
    1. 1.
      ”Development of the indirect‐drive approach to inertial confinement fusion and the target physics basis for ignition and gain” 2267, [35]
    2. 2.
      ”Neoclassical transport of impurities in tokamak plasmas” 1077, [36]
    3. 3.
      ”Measurements of microturbulence in tokamaks and comparisons with theories of turbulence and anomalous transport” 768, [37]
  • Example 3
    User: ”Can you summarize the content of the database?”
    System: The data set appears to be a collection of sentences related to nuclear physics, specifically focusing on topics such as fusion reactions, plasma, and neutron interactions. The sentences describe various relationships between different entities, including deuterium, tritium, neutrons, and other particles, as well as concepts like ignition, implosion, and radiation.
    The data set seems to be a compilation of information from various sources, possibly research papers or scientific articles, and covers a range of topics, including:
    1. Fusion reactions and plasma physics
    2. Neutron interactions and scattering
    3. Ignition and implosion processes
    4. Radiation and heat transfer
    5. Particle physics and nuclear reactions
    The sentences are written in a formal, technical tone, suggesting that the data set is intended for an audience with a background in physics or a related field. Overall, the data set provides a comprehensive overview of various concepts and relationships in nuclear physics, but its purpose and intended use are unclear without further context.
V Conclusion
In this work, we introduced a multi-step approach to automated construction of a KG of nuclear fusion energy. By leveraging the inference power of LLMs, we extracted entities and relationships from 8358 scientific abstracts in the field creating a KG that effectively represents its domain knowledge. This comprehensive representation is not only beneficial for data storage but also serves as a foundation for the development of a KG-RAG machine. Our results highlight the system’s abstraction capabilities over a complex domain knowledge, demonstrating its potential for enhancing information retrieval and knowledge representation in specialized fields.
VI Acknowledgments
This work was supported by the Hartree National Centre for Digital Innovation, a collaboration between STFC and IBM. We also thank the collaboration Fusion Computing Lab between the UKAEA and the STFC Hartree Centre for its partial support to A. Agnello during this work. The author used ChatGPT (OpenAI) to assist with improving the clarity and linguistic expression of the manuscript. All ideas, logical structure, methodology, and conclusions are the result of the author’s own work.
Appendix A Appendix A: list of category types
List of category types along with their corresponding definitions. While these definitions were not used in our analysis and during NER, we include them for clarity.
Entity Type Description
Concept A general idea or theoretical model that is used to understand, describe, or predict phenomena related to fusion processes.
Nuclear Fusion Experimental Facility A specialized research site designed to study and advance nuclear fusion as a viable energy source.
Nuclear Fusion Technique A method or approach designed to achieve and sustain nuclear fusion reactions, for example: Inertial fusion, magnetic confinement.
Nuclear Fusion Device Type Categories of nuclear reactors used in nuclear fusion experiments, for example: tokamak, stellarator.
Nuclear Fusion System Component Key structural or functional parts of a fusion device critical for operation, for example: wall, target, divertor.
Nuclear Fusion System Configuration Physical and operational setups that include detailed descriptions of the devices, systems, and parameters.
Experimental Apparatus Specialized equipment or subsystems used to support fusion experiments, typically external to the main fusion device. For example: vacuum system, cryogenic system, electronic components.
Physical Process Processes governing the interactions and transformations of physical systems.
Physics Entity Quantifiable characteristics that describe the state or behavior of physical systems.
Field Configuration Specific arrangement and intensity of magnetic and/or electric fields used to confine, control, and stabilize the plasma in fusion reactors.
Particle Subatomic particles.
Chemical Element or Compound An element is a pure substance consisting of one type of atom, distinguished by its atomic number. A compound is made of different elements bonded together.
Plasma Property Defines specific attributes or states of plasma in fusion experiments, for example: plasma states and modes (H-mode, edge-localized modes (ELMs)), plasma temperature and density.
Plasma Event Distinct events that occur within plasma during nuclear fusion experiments, for example: disruption, thermal quenching.
Plasma Region Focuses on spatial divisions and areas within a plasma system, for example: core, edge, scrape-off layer.
Plasma Dynamics and Behavior Dynamic processes and interactions that occur within the plasma, for example: plasma transport, turbulent diffusion, magnetohydrodynamic (MHD).
Detection and Monitoring Systems Instruments and tools used to measure and analyze plasma behavior and reactor performance.
Control Systems Systems designed to actively regulate and maintain optimal conditions within the reactor.
Theory and Calculation Theoretical models, frameworks, or computational methods explaining or predicting fusion-related phenomena.
Software and Simulation Computational tools or software for simulating and analyzing fusion systems.
Time Reference A specific temporal reference that defines when an event occurred, such as a particular date, year.
Country and Location Geographical regions (such as countries or cities) or specific places and locations.
Facility or Institution Organizations, laboratories, or research establishments supporting fusion studies.
Person Individuals.
Safety Feature and Regulatory Standard Mechanisms, standards, guidelines ensuring safe operations in nuclear fusion environments.
Database Organized collection of structured information or data.
Scientific Publication and Citation Peer-reviewed written document presenting the results of scientific research or a reference to such a document used to acknowledge its contribution.
Research Field A specialized area of study.